# Tunable Passivity Control for Centralized Multiport Networked Systems

### v1

@@ I-A Literature Review
Centralized networked architectures have become vital across domains, comprising centralized hub nodes interconnecting numerous remote nodes through various network topologies. Example applications encompass distributed multi-agent control systems, complex epidemic models, and complex multilateral telerobotic systems. The topic has been extensively investigated in the area of multilateral telerobotics (eg. Single-Leader Multi-Follower (SLMF) architecture) and several control approaches have been designed for this application, including unmanned exploration, haptic communication, multi-agent control, surgical robot training, and more [a1, a2, a3]. In contrast to bilateral dual-node networked systems, which are typically distinguished by one local node (as the sole leader node) and one remote node (as the follower node), centralized multiport networks combine numerous local node(s) and remote node(s). This technology allows several agents-for example, a human or robotic system or a combination-to interact with each other across various locations through network topology. Example architectures for this type of Centralized Multiport Networked Dynamic (CMND) systems can be found in [a1, a2]. Similar to any networked dynamic architecture, the main challenge is to guarantee and ensure closed-loop stability in the presence of the non-optimum network conditions between each node, besides sensory noises, and actuation faults [i1, i2, i3]. In the last decade, passivity-based controllers have emerged as a solution for both analytical and data-driven control in various applications, including different categories of networked systems [scardovi2010synchronization, 4177251, 1266772, 4287131],. For this, traditionally, wave variable control (WVC)[sun2015wave] and scattering transform control architectures have been proposed and applied in various domains[scardovi2010synchronization], including bilateral telerobotics. More flexible approaches, such as time-domain passivity control (TDPC) and its variety[JHRyu2004, zhouacc2024, paiktim2025, A_Oliver2023MyoPassivity, B_Oliver2023] and variable structure passivity control (VSPC)[Paik2022], have been introduced for dual-node systems. In some relatively recent examples [ex4], an extended TDPC was introduced to the multilateral teleoperation architecture to guarantee stability in the presence of communication time delays. Other examples, such as adaptive controllers based on delay-dependent Lyapunov–Krasovskii functionals, are also proposed [9448498].
@@ II Preliminaries
We present our work for a local node (i.e., hub) communicating with various remote nodes, as can be seen in Fig. 1. An example of this scenario would be in a network telerobotic architecture where various follower robots are controlled remotely via a single leader robot. The sent signal can be the commanded control velocity to the follower robots, and the returned signals would be the reflective force field from each of the remote follower robots of the teleoperation. The signals are sent and received through individual communication channels, which may be different between each node, non-ideal, and have time-varying delays. To ensure the stability and convergence of the entire interconnection, we use the Passivity Theory approach [def1, def3].
@@
where $\xi$ is the inherent Passivity Index (PI) of the local hub[zhou1, zhou2, zhou3, zhouacc2024]. When this value is positive, the hub acts as an inherent dissipation node (i.e., passive system) in the network that dissipates energy with excess of passivity of $\xi$.
@@ III Proposed Method
In this way, we obtain a solution that solves for the adaptive dissipation vector that provides minimal dissipation while guaranteeing the passivity by being subject to the constraint (10). In (11), the weighting matrix ‘$Q$’ is a symmetric positive semi-definite matrix representing the associated penalty for modifying different $\alpha_{i}$. When the associated diagonal element of $Q$ is large, there is a larger penalty for modifying the corresponding $\alpha$, thus effectively reducing the dissipation applied for that node. However, when the associated diagonal element is small, there is a smaller penalty, thus increasing the dissipation applied. The optimization problem can be solved using the Lagrange Multiplier Method [lagrange] as:
@@ IV Simulation Setup
To evaluate and verify the proposed TCoPC framework, we implement the stabilizer in a multilateral teleoperation architecture with a single local hub communicating with three different remote nodes. For telerobotic architectures, the signals sent to the remote nodes are the commanded velocities ‘$v$’, and the received signals are the force feedback ‘$f$’ resulting from the interaction of the follower robots with a remote dynamic environment (see our previous publications for more details on 2-port bilateral telerobotic systems [OSP1, Paik2022]). The simulated parameters of the nodes’ dynamics and the variable communication delays are shown in Table I.
@@ VII Acknowledgements
VIII Appendix
Relation between Strict Passivity and L2 Stability: if the system is an Output Strictly Passive system with EoP of $\xi$, then the system is $L_{2}$ stable with finite $L2$ gain $\leq 1/\xi$ [OSP1, OSP2, hassan]. For a more general form, we prove the relationship between passivity and stability in the continuous time domain. For the OSP system, we have:
$\sum_{k=0}^{N}u^{T}(k)y(k)+E(0)\geq\xi\sum_{k=0}^{N}y^{T}(k)y(k)$ (20)
The condition can be rewritten as:
$\sum_{n=1}^{N}\left(u^{T}[n]y[n]-\xi y^{T}[n]y[n]\right)\geq 0$ (21)
assuming that the initial energy is zero $E(0)=0$. It can be shown that, algebraically
$u^{T}[n]y[n]-\xi y^{T}[n]y[n]$ $=-\frac{1}{2\xi}(u[n]-\xi y[n])^{T}$ (22)
$\quad(u[n]-\xi y[n])+\frac{1}{2\xi}u^{T}[n]u[n]$
$\quad-\frac{\xi}{2}y^{T}[n]y[n]$
so thatas he condition can be rewritten as Let the intermediate terms be defined as
$\mathcal{A}[n]$ $=\frac{1}{2\xi}u^{T}[n]u[n],$
$\mathcal{B}[n]$ $=\frac{\xi}{2}y^{T}[n]y[n].$
The condition can then be rewritten as:
$\sum_{n=1}^{N}\Big(\mathcal{A}[n]-\mathcal{B}[n]\Big)$
$\geq\sum_{n=1}^{N}\Big(-\frac{1}{2\xi}(u[n]-\xi y[n])^{T}(u[n]-\xi y[n])$
$\quad+\mathcal{A}[n]-\mathcal{B}[n]\Big)$ (23)
also, since $\frac{1}{2\xi}(u[n]-\xi y[n])^{T}(u[n]-\xi y[n])$ is negative semi-definite, we know that
$\sum_{n=1}^{N}\Big($ $\frac{1}{2\xi}u^{T}[n]u[n]-\frac{\xi}{2}y^{T}[n]y[n]\Big)$ (24)
$\geq\sum_{n=1}^{N}\Big(-\frac{1}{2\xi}(u[n]-\xi y[n])^{T}(u[n]-\xi y[n])$
$\quad+\frac{1}{2\xi}u^{T}[n]u[n]-\frac{\xi}{2}y^{T}[n]y[n]\Big)$
as a result, combining (VIII) and (24), for the OSP system, we have:
$\sum_{n=1}^{N}\left(\frac{1}{2\xi}u^{T}[n]u[n]-\frac{\xi}{2}y^{T}[n]y[n]\right)\geq 0$ (25)
which we can rewrite as
$\sum_{n=1}^{N}y^{T}[n]y[n]\leq\frac{1}{\xi^{2}}\sum_{n=1}^{N}u^{T}[n]u[n]$ (26)
Therefore, we can prove that all the OSP system is $L_{2}$ stable with a finite $L_{2}$ gain:
$\|y^{T}\|_{L_{2}}\leq\frac{1}{\xi}\|u^{T}\|_{L_{2}}$ (27)

### v2

@@ I-A Literature Review
Centralized networked architectures have become vital across domains, comprising centralized hub nodes interconnecting numerous remote nodes through various network topologies. Example applications encompass distributed multi-agent control systems, complex epidemic models, and complex multilateral telerobotic systems. The topic has been extensively investigated in the area of multilateral telerobotics (eg. Single-Leader Multi-Follower (SLMF) architecture) and several control approaches have been designed for this application, including unmanned exploration, haptic communication, multi-agent control, surgical robot training, and more [19, 6, 15]. In contrast to bilateral dual-node networked systems, which are typically distinguished by one local node (as the sole leader node) and one remote node (as the follower node), centralized multiport networks combine numerous local node(s) and remote node(s). This technology allows several agents-for example, a human or robotic system or a combination-to interact with each other across various locations through network topology. Example architectures for this type of Centralized Multiport Networked Dynamic (CMND) systems can be found in [19, 6]. Similar to any networked dynamic architecture, the main challenge is to guarantee and ensure closed-loop stability in the presence of the non-optimum network conditions between each node, besides sensory noises, and actuation faults [5, 4, 10]. In the last decade, passivity-based controllers have emerged as a solution for both analytical and data-driven control in various applications, including different categories of networked systems [18, 7, 22, 1, 26],. For this, traditionally, wave variable control (WVC)[20] and scattering transform control architectures have been proposed and applied in various domains[18], including bilateral telerobotics. More flexible approaches, such as time-domain passivity control (TDPC) and its variety[17, 25, 13, 12, 11] and variable structure passivity control (VSPC)[14], have been introduced for dual-node systems. In some relatively recent examples [16], an extended TDPC was introduced to the multilateral teleoperation architecture to guarantee stability in the presence of communication time delays. Other examples, such as adaptive controllers based on delay-dependent Lyapunov–Krasovskii functionals, are also proposed [9].
@@ II Preliminaries
We present our work for a local node (i.e., hub) communicating with various remote nodes, as can be seen in Fig. 1. An example of this scenario would be in a network telerobotic architecture where various follower robots are controlled remotely via a single leader robot. The sent signal can be the commanded control velocity to the follower robots, and the returned signals would be the reflective force field from each of the remote follower robots of the teleoperation. The signals are sent and received through individual communication channels, which may be different between each node, non-ideal, and have time-varying delays. To ensure the stability and convergence of the entire interconnection, we use the Passivity Theory approach [21, 8].
@@
where $\xi$ is the inherent Passivity Index (PI) of the local hub[23, 24, 27, 25]. When this value is positive, the hub acts as an inherent dissipation node (i.e., passive system) in the network that dissipates energy with excess of passivity of $\xi$.
@@ III Proposed Method
In this way, we obtain a solution that solves for the adaptive dissipation vector that provides minimal dissipation while guaranteeing the passivity by being subject to the constraint (10). In (11), the weighting matrix ‘$Q$’ is a symmetric positive semi-definite matrix representing the associated penalty for modifying different $\alpha_{i}$. When the associated diagonal element of $Q$ is large, there is a larger penalty for modifying the corresponding $\alpha$, thus effectively reducing the dissipation applied for that node. However, when the associated diagonal element is small, there is a smaller penalty, thus increasing the dissipation applied. The optimization problem can be solved using the Lagrange Multiplier Method [3] as:
@@ IV Simulation Setup
To evaluate and verify the proposed TCoPC framework, we implement the stabilizer in a multilateral teleoperation architecture with a single local hub communicating with three different remote nodes. For telerobotic architectures, the signals sent to the remote nodes are the commanded velocities ‘$v$’, and the received signals are the force feedback ‘$f$’ resulting from the interaction of the follower robots with a remote dynamic environment (see our previous publications for more details on 2-port bilateral telerobotic systems [2, 14]). The simulated parameters of the nodes’ dynamics and the variable communication delays are shown in Table I.